OpenStudy (anonymous):

Find the direct variation equation which passes through (0, 0) and (4, 1).

OpenStudy (anonymous):

We are really finding the equation of the line between these points in y=mx+b form slope, or m = (1-0)/(4-0) or 1.4 to find b, y = mx + b 0 = 0*1/4 + b 0 = b plug these values back in y=mx + b y = 1/4 x + 0 y = 1/4 x